Water enters the condenser at 30oC which has a mass of 10 kg/s. If heat added is 100 kW, determine the outlet temperature in KJ/s.


1
Expert's answer
2021-10-04T13:34:59-0400

For a condenser, we have the following analysis:





Then, by using [eq. (a)],we have to find the temperature of the water leaving the condenser because we can use the information that we have at the start:

"t_{w_{1}}=30°C=303.15\,K
\\ c_w=4.184\frac{kJ}{kg\,K}
\\ m_w=10\frac{kg}{s}
\\ Q=+100\,kW=100\frac{kJ}{s}
\\ Q=h_1-h_2
\\ Q=m_w(h_{w_{2}}-h_{w_{1}})=m_wc_w(t_{w_{2}}-t_{w_{1}})
\\ \implies t_{w_{2}}=t_{w_{1}}+\frac{Q}{m_wc_w}"


After we rearrange, we determine "t_{w_{2}}" or the temperature of the water leaving the condenser:


"t_{w_{2}}=303.15\,K+2.39\,K
\\ \implies t_{w_2}=305.54\,K=32.39\,°C"


With that information we can calculate "h_2=-m_wh_{w_2}" or the 'outlet temperature' using the requested dimensions:


"h_2=-m_wh_{w_2}=-m_w \times c_{w} \times t_{w_2}
\\ h_{2}=-(10\frac{kg}{s})(4.184\frac{kJ}{kg\,K})(305.54\,K)
\\ \implies h_{2}=-12783.79\cfrac{kJ}{s}"


In conclusion, we find the 'outlet temperature' of the condenser as h2 = -12783.79 kJ/s.Reference:

  • Rajput, R. K. (2005). A textbook of engineering thermodynamics. Laxmi Publications.